ISOLDE Technical Coordination Meeting (LS3)
-
-
16:00
→
16:05
General news - Review of minutes 5mSpeakers: Erwin Siesling (CERN), Joachim Vollaire (CERN)
-
16:05
→
16:15
Safety matters 10m
-
16:15
→
16:20
IBDRS project - activities update 5mSpeakers: Ana-Paula Bernardes (CERN), Joachim Vollaire (CERN)
-
16:20
→
16:25
Ventilation primary areas upgrade 5m
Activities to be coordinated with platform installation:
- Gate installation - non conformity to be cleared - finalisation in April with turnstile
- Equipment transfer from 170/3-402 to 197/2-401- 13Feb. Crane placed by route Democrite, - access to the landing 508/170 won't be possible during lifting activities
- Fire detection control rack installation on the platform - 18 Feb
Speakers: Joachim Vollaire (CERN), Melania Averna (CERN) -
16:25
→
16:30
RILIS Laboratory upgrade project - activities update 5mSpeakers: Carlotta Porzio (CERN), Katerina Chrysalidis (CERN), Laura Lambert (CERN), Sebastian Rothe (CERN)
-
16:30
→
16:40
Machne related activities in the experimental hall 10mSpeakers: Erwin Siesling (CERN), Simon Mataguez (CERN)
Refurbishment of the power outlets ongoing (Jean Francois): This morning access to the separator zones was foreseen (with James Devine) but could not take place due to the safety chain being blocked (couldn't take the keys out). Solution (KUKA not running) being looked into.
Check with CV (Xavier Decrand) if the purging of the SRF amplifier circuit (B 199) has been done.
-
16:40
→
16:45
Experiments related activities in the experimental hall 5mSpeakers: Erwin Siesling (CERN), Hanne Heylen (CERN), Oscar Fjeld (CERN)
Cooling water (only demineralized) back in the hall (only user side Low Energy) since Friday 06/02. Checked for leaks/open valves: All ok apart from the purger in the ISCOOL TP circuit with the individual TP valves closed. Closed the ISCOOL cooling water circuit upstream - will need modification anyway in the frame of the RILIS upgrade. No cooling water HIE side. Check chiller for ISS - are they using water?
- PUMA - RC7 needs to be freed up in the coming weeks
-
16:45
→
16:50
Activities in the primary areas - Coactivity management 5mSpeaker: Joachim Vollaire (CERN)
Due to the FUKA Robot being off the conditions for the primary zone safety are missing and the keys for the separator zones could not be taken this monring. Didier Chapuis with Ricardo Martinez Muniz are looking into a solution - mask or take the robot out of the safety chain.
This procedure is now complete (14h30) -
16:50
→
16:55
Activities in the expeirmental hall - Coactivity management 5mSpeakers: Erwin Siesling (CERN), Hanne Heylen (CERN), Oscar Fjeld (CERN)
EPC platform:
Scaffolding for the work on cable trays (and cabling) at the new EPC platfom will be put in place the coming days.Construction area at LA1 for the prefabrication of the platform cable trays - passage kept free (but pay attention).
Racks for the platform coming towards the end of this week (Rustan) - old ventilation handling work taken in account (Melania)
During the arrival of the racks the 508 side entrance will be blocked for a few hours.
MULTIPAC concrete slab next to 508 done last week. Chiller placing foreseen for Monday 09/02 - has not been done yet?
Weekly schedule
- https://isolde.web.cern.ch/isolde-hall-activities-ls3
- Published on Tuesday or Wednesday
- Wed -> Friday (1.5 week)
- on holidays next week
- Filled in before ITC, modified with info from ITC
- Calender for craning from Yann Seraphin is very useful
AOB
- Cleaning:
- Container for cables?
- https://isolde.web.cern.ch/isolde-hall-activities-ls3
-
16:55
→
17:00
Planned visits 5mSpeaker: Hanne Heylen (CERN)
•Visits coordination
•2 slots per week reserved (Wed. afternoon and Friday morning)
•Other slots possible if guides can be found
•Depending on ongoing activities, visits might be restricted to 508 meeting room/part of the ISOLDE hall
Would it be possible to organise an extended visit on Tuesday morning (2-3hours between 09:00-12:00), during which we could set up in a corner to do the recordings in this immersive setting?
-
16:00
→
16:05